OpenStudy (anonymous):

•Solve a system of non-linear equations with a table. •Compare linear, quadratic, and exponential growth functions using a table and a graph. Submit the graph and table. •Identify that exponential functions exceed linear and quadratic functions.

OpenStudy (anonymous):

system of non-linear equations (x^2 + 2y = 4) - (2x + 2y = 1) subtracting the two eqs.(y is eliminated) x^2 - 2x = 3 x^2 - 2x - 3 = 0 Factor this to (x-3)(x+1) = 0 x = +3 and x = -1 the solution to this system

OpenStudy (anonymous):

I dont have idea about table :/